loutalyn5:

A point M on a segment with endpoints X (1, −2) and Y (10, 3) partitions the segment in a 5:1 ratio. Find M. You must show all work to receive credit. (10 points)

MaxyDaKid:

x-coordinate: M = 1 + 5/6*9 = 1 + 7.5 = 8.5 y-coordinate: M = -2 + 5/6*5 = -2 + 4 1/6 = 2 1/6 M = 51/6 13/6 or 8 1/2, 2 1/6

Taybae:

x(1 , -2) ⇒ x₁ = 1 & y₁ = -2 Y(10,3) ⇒ x₂ = 10 & y₂ = 3 m :n = 5 : 1